# A Benchmark of Causal vs. Correlation AI for Predictive Maintenance

### v1

@@ Abstract.
Predictive maintenance in manufacturing environments presents a challenging optimization problem characterized by extreme cost asymmetry, where missed failures incur costs roughly fifty times higher than false alarms. Conventional machine learning approaches typically optimize statistical accuracy metrics that do not reflect this operational reality and cannot reliably distinguish causal relationships from spurious correlations. This study evaluates eight predictive models, ranging from baseline statistical approaches to formal causal inference methods, on a dataset of 10,000 CNC machines with a 3.3 percent failure prevalence. The formal causal inference model (L5) achieved estimated annual cost savings of 1.16 million USD (a 70.2 percent reduction), outperforming the best correlation-based decision tree model (L3) by approximately 80,000 USD per year. The causal model matched the highest observed recall (87.9 percent) while reducing false alarms by 97 percent (from 165 to 5) and attained a precision of 92.1 percent, with a train-test performance gap of only 2.6 percentage points. These results indicate that causal AI methods, when combined with domain knowledge, can yield superior financial outcomes and more interpretable predictions compared to correlation-based approaches in predictive maintenance applications.
@@ 1. Contributions
  • Systematic comparison of causal AI vs. correlation-based machine learning: We evaluate eight models ranging from basic baselines to advanced causal methods on a dataset of 10,000 CNC machines from the UCI Machine Learning Repository [1], providing the first comprehensive benchmark of formal causal inference against traditional ML approaches for predictive maintenance.
  • Business-aligned optimization framework: We demonstrate that models optimized explicitly for business costs (accounting for the 50:1 cost asymmetry between false negatives and false positives) substantially outperform accuracy-optimized approaches, highlighting the critical importance of cost-sensitive learning in industrial applications.
  • Superior financial performance through causal modeling: The formal causal inference model (Model L5) achieved 1,157,500 USD in annual cost savings (70.2% reduction vs. reactive maintenance), outperforming the best correlation-based model (Model L3) by 80,000 USD annually while maintaining identical recall (87.9%).
  • Dramatic reduction in false alarms: The causal model reduced false positives by 97% (from 165 to 5) compared to the best correlation-based approach, achieving 92.1% precision and minimizing operational disruption while maintaining safety through high recall.
  • Enhanced interpretability and actionable insights: Unlike black-box correlation models, the causal approach provides root-cause explanations for each prediction (thermal stress, power overload, or mechanical overstrain), enabling targeted maintenance interventions rather than generic inspections.
  • Robust generalization under distribution shift: Causally-derived features demonstrated stronger generalization (2.6 percentage point train-test gap) compared to raw sensor correlations (4.1 percentage points), consistent with theoretical expectations that causal relationships are invariant under operational changes.
  • Analysis of Bayesian ensemble methods: The PyMC BART model (L7) validates that automatic feature generation can approach causal performance, but without graph-informed constraints it issues substantially more false positives, demonstrating that Bayesian ensembles still benefit from structural causal knowledge.
  • Validation of causal AI value proposition: Results demonstrate that formal causal inference that relies on domain expertise to construct causal graphs delivers superior financial performance, operational efficiency, and interpretability compared to state-of-the-art correlation-based approaches in predictive maintenance applications.
@@ 2. Introduction
Manufacturing faces a maintenance dilemma. Reactive maintenance leads to costly downtime, while overly cautious preventive maintenance wastes resources on healthy equipment. The core challenge is an extreme cost asymmetry across industrial equipment: a missed failure can cost tens of thousands of dollars (unplanned downtime, emergency repairs, production losses), while a false alarm costs only hundreds of dollars (scheduled inspection, standard parts). For the CNC machines analyzed in this study, we use representative costs of 25,000 USD for missed failures and 500 USD for false alarms, reflecting typical manufacturing industry values where the cost ratio between unplanned downtime and preventive maintenance ranges from 20:1 to 100:1 depending on equipment criticality. Table 1 summarizes the four possible prediction outcomes, their associated costs, and business impact.
@@
False Positive (FP) Wrongly predict failure 500 USD Wasted inspection/parts
@@
Standard machine learning models optimize for statistical accuracy metrics such as accuracy, precision, recall, F1-score, and AUC-ROC, which either ignore class imbalance or treat all prediction errors equally. Accuracy is misleading in imbalanced datasets—a model that always predicts “no failure” achieves 96.7% accuracy on our data while providing zero business value. Precision focuses on minimizing false alarms but may miss critical failures, while recall prioritizes catching all failures but may trigger hundreds of unnecessary inspections. F1-score balances precision and recall through their harmonic mean, and AUC-ROC evaluates the model’s ranking ability across all possible thresholds, but neither metric accounts for the actual business costs of different error types. All these metrics fail to capture the 50:1 cost ratio between false negatives and false positives that is typical for manufacturing equipment, where unplanned downtime costs far exceed scheduled maintenance [8]. Furthermore, correlation-based models identify patterns that co-occur with failures but cannot distinguish causation from spurious relationships, leading to poor generalizability and a lack of actionable insights.
Our goal is to determine if Causal AI can deliver superior business value over state-of-the-art, cost-optimized correlation-based ML in a real-world predictive maintenance scenario.
@@ 3. Dataset and Experimental Design
Our analysis utilizes the UCI Machine Learning Repository’s CNC Machine Predictive Maintenance dataset, which contains operational data from 10,000 machines over their production lifecycles. Each machine record includes six sensor measurements collected during normal operation as described in Table 2:
@@
TWF Tool Wear Failure Tool replacement or failure at randomly selected wear time 120
@@
RNF Random Failure 0.1% probability of failure regardless of process parameters (unpredictable) 5
@@ 3.1. Target Definition and Rationale
For this analysis, we focus exclusively on deterministic failures—those failure modes that can theoretically be predicted from sensor measurements and operating conditions. Specifically, our prediction target combines Tool Wear Failure, Heat Dissipation Failure, Power Failure, and Overstrain Failure, while explicitly excluding Random Failure events. Random failures are defined in the dataset as having a fixed 0.1% probability of occurrence regardless of process parameters—they are statistically independent of all sensor readings and represent truly unpredictable component failures (only 5 instances in 10,000 data points). This exclusion reflects a fundamental principle: predictive models should be evaluated on phenomena they can reasonably be expected to predict. Including random failures would unfairly penalize model performance on events that contain no predictive signal by design.
After applying this filter, our dataset contains 9,995 total observations. We employ a stratified 80-20 train-test split, preserving the class distribution in both sets. The test set comprises 2,000 machines, of which 66 experience deterministic failures—a failure rate of 3.3%. This severe class imbalance poses significant challenges for standard machine learning approaches and underscores why accuracy-based metrics are inappropriate for evaluation.
@@
  • Baseline Cost: 1,650,000 USD (reactive maintenance: 66 failures × 25,000 USD)
@@ 3.2.1. Success Metrics
    Total annual cost: The business objective is to minimize the total cost of ownership of the machine
@@
    Recall: Percentage of actual failures caught (which can be critical for safety)
@@
    Precision: Percentage of alarms that are real failures (which can be need to minimize waste)
@@
3.2.2. Correlation-Based Models (L0-L3)
We developed four correlation-based models of increasing sophistication, representing the spectrum of traditional machine learning techniques applied to predictive maintenance.
Model L0: No-Skill Baseline. The no-skill model serves as the minimal performance floor. This model predicts “no failure” for every observation, regardless of sensor readings or operating conditions. While this may seem trivial, it establishes a critical reference point. In highly imbalanced datasets like ours (96.7% non-failure cases), a naive model that always predicts the majority class achieves 96.7% accuracy. This demonstrates why accuracy is a misleading metric for imbalanced problems. The no-skill model catches zero failures, misses all 66 failures in the test set, and incurs the full 1,650,000 USD baseline cost. Any meaningful predictive model must outperform this baseline.
Model L1: Balanced Logistic Regression. The first predictive model applies logistic regression, a fundamental statistical classification technique. Logistic regression models the log-odds of failure as a linear combination of input features: air temperature, process temperature, rotational speed, torque, tool wear, and one-hot encoded machine type. To address class imbalance [10], we employ scikit-learn’s [4] class_weight=‘balanced’ parameter, which automatically adjusts the loss function to penalize misclassification of the minority class (failures) more heavily than the majority class (non-failures).
@@ 3.2.2. Correlation-Based Models (L0-L3)
After training the same logistic regression model as Model L1, we perform a grid search over possible probability thresholds from 0.01 to 0.99 to find the optimal threshold to minimize the total annual cost. For each candidate threshold, we generate predictions on the training set, calculate the resulting confusion matrix, apply our business cost model (5K USD for TP, 500 USD for FP, 25K USD for FN, 0 USD for TN), and compute total cost. The threshold that minimizes total training-set cost is selected as the final decision boundary. This cost-aware optimization explicitly encodes the 50-to-1 cost asymmetry between false negatives and false positives, shifting the decision boundary to favor higher recall even at the expense of increased false alarms.
Note that the threshold is learned on the training set and treated as a learned hyperparameter for test set evaluation.
Model L3: Cost-Aware Decision Tree. While logistic regression models linear relationships, many real-world failure mechanisms involve complex feature interactions. A heat dissipation failure, for example, may only occur when the temperature differential is low AND rotational speed is slow simultaneously - a conjunction that linear models struggle to capture.
@@
3.2.3. Probabilistic and Causal Approaches
To explore whether fundamentally different modeling paradigms can improve upon correlation-based methods, we developed two advanced models incorporating uncertainty quantification and formal causal inference.
Model L4: Bayesian Logistic Regression. While standard logistic regression produces point estimates of failure probability, Bayesian approaches quantify uncertainty in those estimates. Model L4 implements Bayesian logistic regression using PyMC [5], a probabilistic programming framework.
The model architecture mirrors Model L1’s logistic regression structure but replaces point-estimated parameters with probability distributions. Each feature weight is assigned a Normal prior distribution with mean 0 and standard deviation 10, expressing weak prior belief that features have moderate effects. The intercept term receives a Normal(0, 10) prior. Rather than finding a single “best” set of parameters, we use Markov Chain Monte Carlo (MCMC) sampling to explore the posterior distribution of parameters given the data.
Specifically, we run the No-U-Turn Sampler (NUTS) with 500 posterior samples per chain, 300 tuning iterations, and 2 parallel chains, targeting an acceptance rate of 0.85. This produces 1,000 posterior samples representing plausible parameter values. For prediction, we average predicted probabilities across all posterior samples, yielding a probability estimate that incorporates parameter uncertainty.
After obtaining Bayesian predictions, we again optimize the decision threshold for business costs. The key advantage of the Bayesian approach is interpretability: we can quantify uncertainty in predictions and examine posterior distributions of feature weights to assess which sensors provide reliable predictive signal.
Model L5: Causal Model. All previous models (L0-L4) learn correlations in observational data. They identify that certain sensor patterns co-occur with failures but cannot distinguish causation from correlation. Model L5 takes a fundamentally different approach: constructing an explicit causal model and using formal causal inference methods to identify causal effects.
Figure 1. Causal directed acyclic graph (DAG) for CNC machine failures. The graph encodes domain knowledge about physical mechanisms: exogenous variables (sensor readings and machine type) affect endogenous variables (temperature differential, power, and overstrain) which directly cause machine failures.
The causal graph (Figure 1) encodes engineering principles governing CNC machine failure. The graph captures three physical mechanisms:
Thermal failure occurs when insufficient temperature differential (process minus ambient air) prevents heat dissipation, causing component degradation.
Power failure results when mechanical power (torque × rotational speed) exceeds electrical and mechanical capacity.
Overstrain failure reflects cumulative stress where tool wear × torque exceeds material limits.
The graph models these derived variables (temperature differential, power, overstrain) as intermediate nodes mediating sensor effects on failure. This structure reflects CNC physics: sensors combine through physical laws to create stress conditions exceeding thresholds, making causal relationships invariant to operational changes.
The methodology follows a five-step causal inference pipeline
  1. 1)
    Encode domain knowledge as a DAG (Figure 1) with nodes (sensor measurements, machine type, failure) and edges (causal relationships).
  2. 2)
    Define treatment variables: tool wear (replacement timing), torque (operational parameters), and rotational speed (machine settings). Temperature variables are outcomes, not treatments.
  3. 3)
    Apply Pearl’s do-calculus [2] to determine causal effect identifiability. The backdoor criterion [3] identifies adjustment sets blocking confounding paths.
  4. 4)
    Estimate causal effects by fitting linear regressions with identified adjustment sets, obtaining unbiased causal coefficients.
  5. 5)
    Create derived causal features: temperature differential (thermal stress), power (energy dissipation), and overstrain (cumulative mechanical stress) with clear physical interpretations.
Finally, we train a Gradient Boosting Classifier [12] on both the original sensor features and the derived causal features. Gradient boosting builds an ensemble of weak learners (decision trees) that collectively capture complex non-linear relationships. As with previous models, we optimize the decision threshold for business costs.
The critical advantage of this causal approach is that it provides both predictive performance and interpretability. When the model predicts a failure, we can trace the prediction back to specific causal pathways—overstrain exceeding safe limits, thermal stress from inadequate cooling, or power demand exceeding capacity. This transparency enables targeted interventions rather than generic maintenance.
Model L6: Rule-Based Causal Model. We also implement a rule set that mirrors the engineering heuristics described in the dataset documentation: failures are declared when any causal mechanism (tool wear band, insufficient heat dissipation, unsafe power level, or torque-driven overstrain) breaches its learned threshold. Random search tunes those thresholds to minimize cost. The approach provides perfect recall, but, as discussed later, it generates hundreds of precautionary inspections and is therefore best interpreted as an upper bound on achievable recall rather than a deployable policy.
Model L7: PyMC BART. Finally, we train a Bayesian Additive Regression Tree (BART) classifier using the PyMC-BART implementation [14][5]. BART is a non-parametric ensemble that automatically discovers non-linear combinations of features while delivering posterior uncertainty for every prediction. We feed BART the same causal feature set as L5 to test whether a flexible Bayesian learner can match causal performance without the explicit do-calculus pipeline. When we also let BART consume raw sensor readings and automatically generated interaction terms, it rediscovers the same latent constructs (temperature differential, power, overstrain), but the absence of structural constraints increases false positives by 232-3X. This sensitivity analysis confirms the intuition that automatic feature generation alone can approximate causal features, and yet graph-informed regularization is still required to keep operational costs low.
A summary of the modeling techniques is in the table Table 5.

### v2

@@ Abstract.
Predictive maintenance in manufacturing environments presents a challenging optimization problem characterized by extreme cost asymmetry, where missed failures incur costs roughly fifty times higher than false alarms. Conventional machine learning approaches typically optimize statistical accuracy metrics that do not reflect this operational reality and cannot reliably distinguish causal relationships from spurious correlations. This study benchmarks eight predictive models, ranging from baseline statistical approaches to Bayesian structural causal methods, on a dataset of 10,000 CNC machines with a 3.3 percent failure prevalence. While ensemble correlation-based models such as Random Forest (L4) achieve the highest raw cost savings (70.8 percent reduction), the Bayesian Structural Causal Model (L7) delivers competitive financial performance (66.4 percent cost reduction) with an inherent ability of failure attribution, which correlation-based models do not readily provide. The model achieves perfect attribution for HDF, PWF, and OSF failure types. These results suggest that causal methods, when combined with domain knowledge and Bayesian inference, offer a potentially favorable trade-off between predictive performance and operational interpretability in predictive maintenance applications.
@@ 1. Contributions
  • Systematic benchmark of causal vs. correlation-based models: We evaluate eight models ranging from basic baselines to advanced Bayesian structural causal methods on a dataset of 10,000 CNC machines from the UCI Machine Learning Repository [4]. This study provides a comprehensive comparison of formal causal inference against traditional ML approaches for predictive maintenance.
  • Business-aligned optimization framework: We demonstrate that models optimized explicitly for business costs (accounting for the 50:1 cost asymmetry between false negatives and false positives) substantially outperform accuracy-optimized approaches, which highlights the critical importance of cost-sensitive learning in industrial applications.
  • Failure attribution without attribution labels: The Bayesian SCM (L7) attributes each predicted failure to its most likely physical mechanism (TWF, HDF, PWF, or OSF) without requiring failure-type labels during training. This allows for maintenance interventions with a precise focus on the identified root cause.
  • Competitive performance with interpretability: The Bayesian SCM achieves a 66.4% cost reduction, a competitive result compared with the state-of-the-art correlation-based ensemble models. Together with the failure attribution capability, this shows that causal AI can deliver a favorable trade-off between predictive performance and interpretability.
  • Ensemble methods as a performance ceiling: Tree-based ensemble models (Random Forest, AdaBoost, LightGBM) consistently outperform single models including the decision tree, which mirrors the true data-generating process for the synthetic CNC data set. This proves the effectiveness of ensemble methods in reducing variance and preventing overfitting.
@@ 2. Introduction
Manufacturing faces a maintenance dilemma. Reactive maintenance leads to costly downtime, while overly cautious preventive maintenance wastes resources on healthy equipment. The core challenge is an extreme cost asymmetry across industrial equipment: a missed failure can cost tens of thousands of dollars (unplanned downtime, emergency repairs, production losses), while a false alarm may cost only hundreds of dollars (scheduled inspection, standard parts). For the CNC machines analyzed in this study, we use representative costs of 25,000 USD for missed failures and 500 USD for false alarms, reflecting typical manufacturing industry values where the cost ratio between unplanned downtime and preventive maintenance ranges from 20:1 to 100:1 depending on equipment criticality. Table 1 summarizes the four possible prediction outcomes, their associated costs, and business impact.
@@
False Positive (FP) Incorrectly predict failure 500 USD Wasted inspection/parts
@@
Standard machine learning models optimize for statistical accuracy metrics such as accuracy, precision, recall, F1-score, and AUC-ROC, which either ignore class imbalance or treat all prediction errors equally. Accuracy is misleading in imbalanced datasets. In fact, a model that always predicts “no failure” achieves 96.7% accuracy on our data while providing no business value. Precision focuses on minimizing false alarms but may miss critical failures, while recall prioritizes catching all failures but may trigger hundreds of unnecessary inspections. F1-score balances precision and recall through their harmonic mean, and AUC-ROC evaluates the model’s ranking ability across all possible thresholds, but neither metric accounts for the actual business costs of different error types. All these metrics fail to capture the 50:1 cost ratio between false negatives and false positives that is typical for manufacturing equipment, where unplanned downtime costs far exceed scheduled maintenance [8]. Furthermore, correlation-based models identify patterns that co-occur with failures but cannot distinguish causation from spurious relationships. This limits their utility to a binary alarm with no actionable diagnosis.
Our goal is to determine if Causal AI can deliver competitive business value relative to state-of-the-art correlation-based ML while also offering reliable insights into the failure root causes in predictive maintenance scenarios.
@@ 3. Dataset and Experimental Design
Our analysis utilizes the UCI Machine Learning Repository’s CNC Machine Predictive Maintenance dataset, which is a synthetic dataset designed in [4] to reflect real-world predictive maintenance scenarios. The dataset contains 10,000 machines and for each machine six sensor measurements as described in Table 2:
@@
TWF Tool Wear Failure Tool replacement or failure at randomly selected wear time 46
@@
RNF Random Failure 0.1% probability of failure regardless of process parameters (unpredictable) 19
@@ 3.1. Target Definition and Rationale
For this analysis, we focus exclusively on predictable failures—those failure modes that can theoretically be predicted from sensor measurements and operating conditions. Specifically, our prediction target combines Tool Wear Failure, Heat Dissipation Failure, Power Failure, and Overstrain Failure, while explicitly excluding Random Failure events. Random failures are defined in the dataset as having a fixed 0.1% probability of occurrence regardless of process parameters. They are statistically independent of all sensor readings and represent truly unpredictable noise (only 19 instances in 10,000 data points). It is worth noting that while TWF is not excluded, it is largely nondeterministic. In fact, the data generation process of TWF contains an irreducible stochastic component. Readers may refer to [5] or Section 4.1.1 for more details.
After applying this filter, our dataset contains 9,981 total observations. We employ a stratified 80-20 train-test split that preserves the class distribution in both sets. The test set comprises 2,000 machines among which 66 experience deterministic failures (3.3% failure rate). This severe class imbalance poses significant challenges for standard machine learning approaches and underscores why accuracy-based metrics are inappropriate for evaluation.
@@
  • Baseline Cost: 1,650,000 USD (reactive maintenance: 66 failures $\times$ 25,000 USD)
@@ 3.2.1. Success Metrics
    Total annual cost: The business objective is to minimize the total cost of machine maintenance
@@
    Recall: Percentage of actual failures caught
@@
    Precision: Percentage of alarms that are real failures
@@
3.2.2. Correlation-Based Models (L0–L3)
We start by investigating four correlation-based models, ranging from a no-skill baseline through logistic regression variants to a decision tree.
Model L0: No-Skill Baseline. The no-skill model serves as the minimal performance floor. This model predicts “no failure” for every observation, regardless of sensor readings or operating conditions. While this may seem trivial, it establishes a critical reference point. In highly imbalanced datasets like ours (96.7% non-failure cases), a naive model that always predicts the majority class achieves high accuracy. This demonstrates why accuracy is a misleading metric for imbalanced problems. The no-skill model catches zero failures, misses all 66 failures in the test set, and incurs the full 1,650,000 USD baseline cost. Any meaningful predictive model must outperform this baseline.
Model L1: Balanced Logistic Regression. The first predictive model applies logistic regression, which is a fundamental classification technique. Logistic regression models the log-odds of failure as a linear combination of input features: air temperature, process temperature, rotational speed, torque, tool wear, and one-hot encoded machine type. To address class imbalance [2], we employ scikit-learn’s [6] class_weight=‘balanced’ parameter, which automatically adjusts the loss function to penalize misclassification of the minority class (failures) more heavily than the majority class (non-failures).
@@ 3.2.2. Correlation-Based Models (L0–L3)
After training the same logistic regression model as Model L1, we perform a grid search over possible probability thresholds from 0.01 to 0.99 to find the optimal threshold to minimize the total annual cost. For each candidate threshold, we generate predictions on the training set, calculate the resulting confusion matrix, apply our business cost model (5K USD for TP, 500 USD for FP, 25K USD for FN, 0 USD for TN), and compute total cost. The threshold that minimizes total training-set cost is selected as the final decision boundary. This cost-aware optimization explicitly encodes the 50-to-1 cost asymmetry between false negatives and false positives. It encourages the model to shift the decision boundary to favor higher recall even at the expense of increased false alarms. Note that the threshold is learned on the training set and treated as a learned hyperparameter for test set evaluation.
Model L3: Cost-Aware Decision Tree. While logistic regression models linear relationships, many real-world failure mechanisms involve complex feature interactions. A heat dissipation failure, for example, may only occur when the temperature differential is low AND rotational speed is slow simultaneously. This presents a conjunction that linear models struggle to capture.
@@
3.2.3. Ensemble Models (L4–L6)
Model L4: Random Forest. While the decision tree model is akin to the true data generating process of this dataset, in practice a single tree often suffers from high variance. A small change in training data can yield drastically different tree structures. Thus we consider the Random Forest model, which addresses this limitation via ensembling. A Random Forest model builds multiple decision trees and aggregates their results by averaging their predictions. This aggregation method reduces variance compared to using a single model, and helps prevent overfitting. In fact, during training we employed the following configuration:
  1. (1)
    Bootstrap sampling: Each tree is trained on a bootstrap sample, meaning that we draw $N$ training samples with replacement from the original $N$ samples. Each tree sees approximately 63% unique samples.
  2. (2)
    Random Feature Selection: We randomly select $\sqrt{7}\approx 3$ features as candidates when splitting a node. This prevents strong features from dominating every tree’s upper splits and encourages discovering diverse patterns.
  3. (3)
    Regularization Constraints: We require at least 10 samples to split a node. This helps prevent overfitting.
The Random Forest model is implemented using the scikit-learn library. After obtaining ensemble probability predictions, we apply cost-aware threshold optimization as in Models L2--L3 to minimize total business cost.
Model L5: AdaBoost. Now we turn to boosting algorithms. We first consider AdaBoost, which is a foundational boosting algorithm that demonstrates the strength of combined weak learners. AdaBoost trains the weak learners on a weighted version of the training data, where the weight on incorrectly classified samples increases exponentially after each iteration.
We implement AdaBoost using scikit-learn’s AdaBoostClassifier with decision stumps (depth-1 trees) as base learners. We train 50 decision stumps with learning rate 1.0. Each stump makes a binary decision based on one feature. The ensemble prediction is a weighted average over all stumps.
Finally, we apply the cost-aware threshold optimization as in Models L2--L3.
Model L6: Light Gradient Boosting Machine. As the final model in the ensemble class, we consider LightGBM, which represents the modern evolution of gradient boosting algorithms. It greatly reduces training time compared to earlier gradient boosting algorithms (e.g. XGBoost, CatBoost) while maintaining or improving accuracy. For more details regarding the innovations of LightGBM, the reader may refer to [3].
We implement LightGBM using the lightgbm Python library with 100 boosting iterations, maximum tree depth of 3, learning rate of 0.1, and 10 maximum leaves per tree. Cost-aware threshold optimization is applied at the end of the pipeline.
3.2.4. Causal Model
Finally, we consider a causal model that utilizes domain-specific failure mechanisms. This structure allows for both prediction and failure attribution, i.e. telling which failure mode (TWF, HDF, PWF, or OSF) is responsible.
Model L7: Bayesian Structural Causal Model (SCM). Unlike correlation models that learn $P(Y|X)$ by fitting weights to all available features, this model encodes the specific physical mechanisms of failure into a structural graph. We explicitly model four independent failure mechanisms that combine via a logical “Noisy-OR” gate to produce the final machine failure.
We define the system using a set of structural equations implemented in PyMC [7]. We standardized all input features $X$. In the following, we let $\sigma(z)=1/(1+e^{-z})$ be the sigmoid function.
1. Tool Wear Failure (TWF) Mechanism: This mechanism depends exclusively on the tool wear duration. We model the probability $p_{twf}$ as a logistic function of tool wear ($x_{tw}$), excluding all other variables to enforce causal independence:
(1) $p_{twf}=\sigma(\alpha_{twf}+\beta_{twf}\cdot x_{tw}).$
2. Heat Dissipation Failure (HDF) Mechanism: Physical heat failure occurs when the heat cannot be dissipated, which happens when the temperature difference ($x_{\Delta T}$) is low AND rotational speed ($\omega$) is low. We model this using an explicit interaction term:
(2) $p_{hdf}=\sigma(\alpha_{hdf}+\beta_{temp}x_{\Delta T}+\beta_{rot}\omega+\beta_{int}(x_{\Delta T}\cdot\omega)).$
3. Power Failure (PWF) Mechanism: Power failure occurs if the power ($P=\tau\cdot\omega$) is too low or too high. To capture this non-monotonic U-shaped relationship, we introduce a quadratic term $P^{2}$:
(3) $p_{pwf}=\sigma(\alpha_{pwf}+\beta_{p}P+\beta_{p^{2}}P^{2}).$
4. Overstrain Failure (OSF) Mechanism: Overstrain depends on the product of Tool Wear and Torque ($S=x_{tw}\cdot\tau$). The failure threshold depends on the machine quality type ($L,M,H$). We model this using dummy variables for Low ($\mathbb{I}_{L}$) and Medium ($\mathbb{I}_{M}$) types:
(4) $p_{osf}=\sigma(\alpha_{osf}+\beta_{osf}S+\beta_{L}\mathbb{I}_{L}+\beta_{M}\mathbb{I}_{M}).$
5. System Failure (Noisy-OR): Finally, we combine these independent probabilities using a Noisy-OR gate. This logic implies that the machine survives only if it survives all individual failure mechanisms:
(5) $P(\text{Failure})=1-\left[(1-p_{twf})(1-p_{hdf})(1-p_{pwf})(1-p_{osf})\right].$
We place weakly informative Normal priors on all $\alpha$ and $\beta$ parameters (e.g., $\mathcal{N}(0,1)$) and use NUTS sampling to estimate the posterior distributions.
A summary of the modeling techniques is in Table 5.